OpenStudy (anonymous):

Graph the first six terms of a sequence where a1 = -10 and d = 3.

OpenStudy (anonymous):

A)

OpenStudy (anonymous):

B)

OpenStudy (anonymous):

C)

OpenStudy (anonymous):

D)

OpenStudy (anonymous):

@iGreen @Michiganboy21

OpenStudy (igreen):

Arithmetic Sequence? Or Geometric?

OpenStudy (anonymous):

graphing sequence @iGreen

OpenStudy (igreen):

-10 + 3 = -7 -7 + 3 = -4 -4 + 3 = -1 -1 + 3 = 2

OpenStudy (igreen):

2 + 3 = 5 5 + 3 = 8

OpenStudy (igreen):

Those are the six terms.

OpenStudy (anonymous):

where did you get those numbers

OpenStudy (igreen):

Start with \(a_1\), which is -10. Common difference is 3. So you keep adding 3

OpenStudy (igreen):

Your answer is A.

OpenStudy (igreen):

The solutions to those 6 problems I did before are the y values. The x values are 1, 2, 3, 4, 5, 6.